Revenue and Elasticity (6 points)

Amazon.com is trying to increase its total revenue. It explores a 10% discount on its products. Based on internal research, amazon knows that its customers can be divided into two distinct groups (Prime and General). The following table shows how the two groups respond to the discount (amounts are sales/week). Assume that all products sold cost the same (production and price).

General Customers

Amazon Prime Customers

Volume of sales before the 10% discount

1.50 million

1.55 million

Volume of sales after the 10% discount

1.70 million

1.65 million

a) Calculate the price elasticity of demand for General Customers and for Prime Customers.

b) Explain how the discount will affect total revenue from each group.

c) In order to maximize revenue, should discounts be offered to General Customers only, to Prime only, to neither, or to both? Why?
Answer #1

Ans

General=1.70-1.50/1.50 ehole divided by 10%=4

Prime=1.65-1.55/1.55 whole divided by 10% or 0.1=0.645

2 revenue will rise in case of general because demand is elastic (>1) and decrease in case of prime because demand is inelastic

3 To general customers only because discount increases revenue only in their case